OpenStudy (anonymous):

find the equation of the line with given slope and y-intercept. m=-3, y-int: (0,3); slope-intercept form

OpenStudy (anonymous):

Do you know the equation for slope-intercept form?

OpenStudy (anonymous):

@Xavier no

OpenStudy (anonymous):

Given : m/slope = -3, y-intercept : (0,3) slope intercept form : y=mx+b therefore. y=(-3)x+3 y=-3x+3 3x+y-3=0